pH of a solution produced when an aqueous solution of pH 6 is mixed with an equal volume of aqueous solution of pH 3 is about__.
 
Fill in the blank

Grade:12

2 Answers

Khimraj
6 years ago
Let volume be V then new H+ concept will be [(10^-6)V + (10^-3)V]/2V so it dual to approx 10^-3/2 so pH of resultant sol is 3.3.Hope it clears. If u have doubts then please clearify. And if u like my answer then please approve answer
Shrimay kher
6 years ago
pH = 6 
-log [H+] = 6
[H+]1 =10-6
pH = 3 
-log [H+] =3
[H+]2 =10-3
[H+]mix= [H+]1V1 + [H+]2V/ V1+V2
V1=V2=V
[H+]mix = (10-6+10-3)V/2V
           = (10-6+10-3)/2
 10-6 being small neglected
           = 0.5 x10-3
 [H+]mix = 5 x 10-4
 pHmix = – log[H+]mix
              = 4 – log5
          = 4 – 0.7            pHmix  = 3.3